OpenStudy (needhelpwithk12):

HELP PLEASE WILL GIVE MEDAL TO PERSON THAT ANSWERS 5 QUESTIONS FOR ME!!! Solve for x. Express the answer in simplest form. 1/7=3/4X-1

OpenStudy (needhelpwithk12):

@likeabossssssss

OpenStudy (needhelpwithk12):

hello?anyone???

OpenStudy (palwandew):

wait a minute.

OpenStudy (needhelpwithk12):

ok

OpenStudy (needhelpwithk12):

this is what I got so far. 1/7=12x-3

OpenStudy (palwandew):

First cross multiply.

OpenStudy (needhelpwithk12):

I already have i think?

OpenStudy (needhelpwithk12):

@ILovePuppiesLol

OpenStudy (greenbannanabob):

1/7=3/4X-1 Multiply both sides by 7 1=21/4x-7 Multiply both sides by 4x 4x=21-28x Add 28x to both sides 32x=21 Divide 21 by 32 x=???

OpenStudy (greenbannanabob):

I first got rid of all the fractions cause they annoy me(might seem like more work but it makes it more organized) and then isolated x.

OpenStudy (needhelpwithk12):

i need a fraction.

OpenStudy (greenbannanabob):

Your answer will come out as a fraction

OpenStudy (needhelpwithk12):

oh it will?

OpenStudy (palwandew):

x=22/4 should be your answer

OpenStudy (needhelpwithk12):

what I got when I divided 21 by 32 was 0.65625

OpenStudy (needhelpwithk12):

ok next one

OpenStudy (needhelpwithk12):

Solve for x.